how do you evaluate the logarithm 3 log(base2)of2 - log(base2)of4?
Answer by lwsshak3(6490)

how do you evaluate the logarithm 3 log(base2)of2 - log(base2)of4?
3 log(base2)of2 - log(base2)of4
log2(2^3)-log2(4) (power rule)
log2(2^3)/4 (division rule)
log2(8/4)
log2(2)=1 (The logarithm of its base is always equal to 1, that is, the base,2, must be raised to what number(logarithm=1) to equal 2.

ans: (3 log(base2)of2 - log(base2)of4) = 1
check: 3log2(2)=3
log2(4)=log2(2^2)=2
3-2=1